Recent Log Entries

COMMENT ON Injector Upgrade - Phase 1 Installation (Gun to Chopper)

Log Number: 3853428Logbook(s): Body: 

DecaRad boxes 1&2, cables and GM tube heads have been
removed from the beam line. Locations have been documented
and materials are stored in cabinet marked INJ Lab 4.
Port Server injdigips was connected to switch opsin01sw2 #20.
Port 1 DecaRad1
Port 2 DecaRad2
Port 3
Port 4 DTM 151 Tesla Meter

COMMENT ON PSS Injector Interface Chassis removal

Log Number: 3853426Logbook(s): Body: 

Lock Out and Tag Out of hazards in the Injector segment will be handled under existing ATLIs for this purpose. It covers Gun HVPS and RF zones.
Injector segment PSS system will be powered down for the upgrade of the system.
PSS Interface chassis will be disconnected and removed for the upgrade.
Separate ATLis is used to track the upgrade work.
Task discussed with owner and approved.

VWien soak ended

Log Number: 3853422Logbook(s): Entry Makers: gramesBody: 

The HWien appears to induce a very modest ~0.5 nA vacuum signal increase at VIP0I01 when I tested it last week, with onset +/- 4 kV (note 90 deg is +/- 11 kV), and did NOT scale with voltage. Interestingly, +/- 4 kV is close to the ~27 deg we ran at for months. So, I soaked the VWien at +/- 11 kV (in the POS configuration over the weekend). The vacuum activity I observed was not apparent, see Fig. 1.

So, we're in good shape. Both Wiens are turned OFF in software now, we'll unpower and safe the PS and disconnect HV cables.

COMMENT ON Injector Upgrade - Phase 1 Installation (Gun to Chopper)

Log Number: 3853419Logbook(s): Body: 

Day #6, Week2 Day1, Post-830AM Toolbox Meeting, Plan for the
Day:

* Power down Wein system; InjSB + tunnel (Scott/Joe)
* Power down and remove DecaRad system; tunnel (Scott)
* Return gun valve config to "normal"; tunnel (Marcy)
* Dispose of grey mineral oil tank in coordination w/ESH&Q'
InjSB (Carlos/Bubba)
* PSS Kickers and associated chassis removal; InjSB + tunnel
(SSG)
* Laser table disassembly; tunnel (Shukui/Scott)

Other work possibly occurring upstairs in InjSB:
* Harp upgrades @IN03B07 (Dela Cruz)
* Chopper amplifier upgrades, IN01 racks (Farrish)

https://wiki.jlab.org/ciswiki/index

ABIL tag #3789 on Faraday Cup #1 has been removed

Log Number: 3853414Logbook(s): Body: The following entry was removed from the Accelerator Bypassed-Interlocks Log (ABIL) Tag # 3789 Device: Faraday Cup #1 Function: Insert Location: IFY0I06T Purpose: FC1 bellows has repaired leak but is not robust enough to insert.

2L06,07,20,21 LL to 20%

Log Number: 3853349Logbook(s): Body: 

Need more room for gas expansion after CHL1 shutdown/North Linac warmup/Hall A warmup. Taking a few South Linac cryomodules to 20% and shipping the inventory away in the tanker truck. We should be able to fill them back up end of next week.

Pages